OpenStudy (anonymous):

what is the fraction for 0.207

OpenStudy (anonymous):

\[.207 = \frac{.207}{1} = \frac{2.07}{10}= \frac{20.7}{100}= \frac{207}{1000}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Do you understand the process by which you solve these?

OpenStudy (anonymous):

yes i understand